# Flat or crumpled: states of active symmetric membranes

### v2

@@ Appendix A Derivation of the hydrodynamic equations
to the leading order in $h$-fluctuations. In deriving the above equations, we have used chaikin
@@
Appendix B Nonlinear terms
We now systematically derive the nonlinear terms which may be added to the dynamical equations for $h,u_{\parallel}$ and $u_{\perp}$. There are several sources of possible nonlinearities. We first consider those which are spatially local. These can come from (i) $v_{\text{perm}}$ as given in the main text, which has both active and equilibrium contributions, (ii) $(\delta{\cal F}/\delta h)$ that contributes to equilibrium relaxation, and (iii) active stress $\sigma_{ij}^{a}$. To proceed further, we start from the form of the nonlinear stress nelson :
@@ Appendix B Nonlinear terms
neglecting the the contributions quadratic in $\bf u$ nelson . This gives
@@
Appendix C Solutions of the bulk flow fields
We now solve for $v_{z}^{\prime}({\bf q},z,t)$. We note that
$v_{z}^{\prime}({\bf q},z=0,t)=A_{1}({\bf q},t)=A_{2}({\bf q},t).$ (56)
Next, by using the results obtained in Section A, we find
Solving
$B_{1}$ $=$ $qv_{z}^{\prime}({\bf q},z=0,t)-iqv_{\parallel}({\bf q},t),$ (58)
$B_{2}$ $=$ $-qv_{z}^{\prime}({\bf q},z=0,t)-iqv_{\parallel}({\bf q},t).$ (59)
We thus obtain

### v3

@@ I Derivation of the hydrodynamic equations
to the leading order in $h$-fluctuations. In deriving the above equations, we have used chaikin-sm
@@
II Solutions of the bulk flow fields
We now solve for $v_{z}^{\prime}({\bf q},z,t)$. We note that
$v_{z}^{\prime}({\bf q},z=0,t)=A_{1}({\bf q},t)=A_{2}({\bf q},t).$ (50)
Next, by using the results obtained in Section I, we find
Solving
$B_{1}$ $=$ $qv_{z}^{\prime}({\bf q},z=0,t)-iqv_{\parallel}({\bf q},t),$ (52)
$B_{2}$ $=$ $-qv_{z}^{\prime}({\bf q},z=0,t)-iqv_{\parallel}({\bf q},t).$ (53)
We thus obtain
III Nonlinear terms
We now systematically derive the nonlinear terms which may be added to the dynamical equations for $h,u_{\parallel}$ and $u_{\perp}$. There are several sources of possible nonlinearities. We first consider those which are spatially local. These can come from (i) $v_{\text{perm}}$ as given in the main text, which has both active and equilibrium contributions, (ii) $(\delta{\cal F}/\delta h)$ that contributes to equilibrium relaxation, and (iii) active stress $\sigma_{ij}^{a}$. To proceed further, we start from the form of the nonlinear stress nelson-sm :
@@ III Nonlinear terms
neglecting the the contributions quadratic in $\bf u$ nelson-sm . This gives